A piece of titanium metal with a mass of 20.8 g is

heated in boiling water to 99.5°C and then dropped

into a coffee cup calorimeter containing 75.0 g of

water at 21.7°C. When thermal equilibrium is reached,

the final temperature is 24.3°C. Calculate the specific

heat capacity of titanium.


1
Expert's answer
2021-08-11T05:22:37-0400

Solution:

1) This problem can be summarized thusly:

-qlost by titanium = qgained by water

 

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J g-1 °C-1)

Tf = final temperature

Ti = initial temperature

The specific heat capacity of water is 4.20 J g-1 °C-1

 

2) Therefore:

qlost by titanium = (20.8 g) × Ctitanium × (24.3 - 99.5)°C

qlost by titanium = -1564.16 × Ctitanium

 

qgained by water = (75.0 g) × (4.20 J g-1 °C-1) × (24.3 - 21.7)°C

qgained by water = 819

 

-qlost by titanium = qgained by water

-(-1564.16 × Ctitanium) = 819

Ctitanium = (819) / (1564.16) = 0.5236 = 0.524

Ctitanium = 0.524 J g-1 °C-1


Answer: The specific heat capacity of titanium is 0.524 J g-1 °C-1